OpenStudy (anonymous):

PLZ HELPP!!! A cyclist rides her bike at a speed of 20 miles per hour. What is this speed in kilometers per hour? How many kilometers will the cyclist travel in 3 hours? In your computations, assume that 1 mile is equal to 1.6 kilometers. Do not round your answers.

OpenStudy (anonymous):

@rational

OpenStudy (rational):

1 mile = 1.6 kilometers 20 miles = ? kilometers

OpenStudy (anonymous):

@rational huh?

OpenStudy (anonymous):

@rational i dont understnd i dont know

OpenStudy (rational):

1 mile is same as 1.6 kilometers so 20 miles is same as how many kilometers ?

OpenStudy (anonymous):

@rational i still no understand

OpenStudy (anonymous):

20miles=12.5kilimeters

OpenStudy (rational):

12,5 kilometers is wrong

OpenStudy (rational):

idk how to explain u better

OpenStudy (anonymous):

@rational umm is it 20.6?

OpenStudy (rational):

1 mile = 1.6 kilometer so 20 miles = 20 * 1.6 kilometers = 32 kilometers

OpenStudy (anonymous):

@rational wait wait

OpenStudy (anonymous):

@rational i also need speed in km/h

OpenStudy (anonymous):

@rational how i do?

OpenStudy (rational):

yeah speed itself is `32 kilometers per hour`

OpenStudy (rational):

multiply that by 3 to get the distance travelled in 3 hours

OpenStudy (anonymous):

and distance in 3 hours how much? @rational

OpenStudy (anonymous):

ohh ok

OpenStudy (rational):

distance travelled in 3 hours = 3*32 = 96 kilometers

OpenStudy (anonymous):

@rational thank you

OpenStudy (rational):

you're welcome!